# WorldForge: Unlocking Emergent 3D/4D Generation in Video Diffusion Model via Training-Free Guidance

### v1

@@ Abstract
Recent video diffusion models demonstrate strong potential in spatial intelligence tasks due to their rich latent world priors. However, this potential is hindered by their limited controllability and geometric inconsistency, creating a gap between their strong priors and their practical use in 3D/4D tasks. As a result, current approaches often rely on retraining or fine-tuning, which risks degrading pretrained knowledge and incurs high computational costs. To address this, we propose WorldForge, a training-free, inference-time framework composed of three tightly coupled modules. Intra-Step Recursive Refinement introduces a recursive refinement mechanism during inference, which repeatedly optimizes network predictions within each denoising step to enable precise trajectory injection. Flow-Gated Latent Fusion leverages optical flow similarity to decouple motion from appearance in the latent space and selectively inject trajectory guidance into motion-related channels. Dual-Path Self-Corrective Guidance compares guided and unguided denoising paths to adaptively correct trajectory drift caused by noisy or misaligned structural signals. Together, these components inject fine-grained, trajectory-aligned guidance without training, achieving both accurate motion control and photorealistic content generation. Extensive experiments across diverse benchmarks validate our method’s superiority in realism, trajectory consistency, and visual fidelity. This work introduces a novel plug-and-play paradigm for controllable video synthesis, offering a new perspective on leveraging generative priors for spatial intelligence.
Figure 1: We present WorldForge, a fully training-free framework leveraging a pre-trained video diffusion model for various 3D/4D tasks, such as monocular 3D scene generation (up) and dynamic 4D scene re-rendering (down), enabling precise camera trajectory control and high-quality outputs.
@@ 1 Introduction
Recent advances in generative modeling, particularly video diffusion models (VDM) (Blattmann et al., 2023; Wan et al., 2025; Yang et al., 2024; Google DeepMind, 2025), have greatly expanded the scope and capabilities of spatial intelligence (Cao et al., 2025) tasks such as 3D and 4D scene understanding (Bahmani et al., 2025a; b), reconstruction (Wang et al., 2025a; Wu et al., 2025; Shi et al., 2024), and generation (Yu et al., 2024c; 2025). Trained on massive, diverse video datasets, these models inherently encode rich spatiotemporal priors that capture structural, temporal, and motion-related patterns. Leveraging such priors offers significant advantages for achieving realistic and coherent spatial transformations, enabling applications in novel view synthesis (You et al., 2025; Xiao et al., 2025), panoramic video generation (Wang et al., 2024b; Ma et al., 2024a), 3D scene generation (Liu et al., 2024; Liang et al., 2025; Wang et al., 2024c), and dynamic scene reconstruction (Bai et al., 2025a; Yu et al., 2025; Van Hoorick et al., 2024). Moreover, VDMs are increasingly used to construct “world models” (Bar et al., 2025; Duan et al., 2025; Bruce et al., 2024), understood as structured internal representations of physical environments that support predictive reasoning, planning, and decision-making in embodied AI systems.
@@
Despite their strong generative capabilities, current video diffusion models face fundamental limitations, including limited controllability, spatial–temporal consistency, and geometric fidelity, particularly when applied to 3D or 4D tasks (Wang et al., 2024c; He et al., 2024; Ling et al., 2024; Xing et al., 2024). While they can be loosely guided with text prompts or reference frames (Kong et al., 2024; Wan et al., 2025), they generally lack the ability to follow precise user-defined motion constraints, such as a specified 6-DoF camera trajectory or object pose evolution (Hu, 2024; Ma et al., 2024b). This lack of trajectory-level control is especially problematic in tasks requiring spatial consistency across views, such as novel view synthesis or free-viewpoint rendering. Furthermore, these models often entangle scene and camera motion, making it difficult to separate object dynamics from viewpoint changes (Yu et al., 2024c; Liu et al., 2024). Consequently, attempts to synthesize videos with fixed scenes or smooth camera paths often result in unintended object deformations or scene instability. These limitations hinder their applicability in domains requiring structured spatial reasoning or controllable video generation.
@@
To handle these limitations, prior works (Jeong et al., 2025; Ren et al., 2025; Yu et al., 2025; Zhang et al., 2025) have explored two main directions. The first involves training or fine-tuning the generative backbone on multiview or motion-conditioned data, often with explicit modules to encode trajectory information (Bai et al., 2025a; Xiao et al., 2024; Bai et al., 2025b). While this approach can improve alignment and control, it is computationally costly, may generalize poorly to diverse scenes, and risks degrading the model’s pretrained priors during fine-tuning. The second line of work adopts a “warping-and-repainting” strategy (Ma et al., 2025b; Liu et al., 2025; Ma et al., 2025a; You et al., 2025), in which input frames are lifted into a partial 3D representation (e.g., via depth estimation (Piccinelli et al., 2024; Yin et al., 2023)), re-projected along a user-defined camera path, and then refined by a generative model to fill missing regions. Although more flexible, such methods face notable robustness limitations, as pretrained models are not designed to process out-of-distribution (OOD) (Yu et al., 2024a) inputs such as warped or disoccluded images. As a result, they often produce artifacts and missing structures (e.g., suspended components or fragmented geometry). Moreover, their bias toward dynamic training data leads to hallucinated motion even in static scenes, undermining view consistency. In summary, the combined effects of OOD inputs and dynamic-data bias make it challenging to balance fine-grained controllability with generation quality and generalization, which remains an open problem.
@@
To address this challenge, we aim to inject precise control into VDMs while preserving their valuable priors. For this purpose, we propose a general inference-time guidance paradigm that leverages the rich priors of large-scale VDMs (Blattmann et al., 2023; Wan et al., 2025) in spatial intelligence tasks, such as geometry-aware 3D scene generation and video trajectory control. Our method adopts a warping-and-repainting pipeline, in which input frames are warped along a reference trajectory and then used as conditional inputs in the repainting stage. Building on this, we develop a unified, training-free framework composed of three complementary mechanisms, each designed to address a specific challenge in trajectory-controlled generation.
@@
First, to ensure the generated motion follows the target trajectory derived from depth-based rendering (Wang et al., 2025b; Piccinelli et al., 2024), we introduce Intra-Step Recursive Refinement (IRR). It embeds a micro-scale predict–correct loop within each denoising step: before the next timestep, predicted content in observed regions is replaced with the corresponding ground-truth observations. This incremental correction allows trajectory control signals to be injected at every step, enabling fine-grained, stepwise guidance that keeps the motion aligned with the target trajectory.
@@
Second, we observe that different channels of the VAE-encoded (Kingma & Welling, 2013; Foti et al., 2022) latent representation encode different information, with some channels specializing in appearance and others in motion. Directly overwriting all channels when injecting trajectory signals can inadvertently degrade fine-grained visual details. To address this, we propose Flow-Gated Latent Fusion (FLF), which leverages optical-flow similarity to selectively and dynamically inject trajectory information only into motion-relevant channels, while leaving appearance-relevant channels unmodified. This selective modulation effectively decouples appearance from motion, allowing for precise viewpoint manipulation while preserving content fidelity.
@@
Finally, while warping-based rendering effectively enforces user-defined trajectories in view synthesis, it inevitably introduces noise stemming from imperfect depth estimation, occlusions, and scene misalignments. These imperfections often lead to visual artifacts such as ghosting, structural distortions, and degraded temporal coherence (You et al., 2025). To better balance trajectory adherence and generation fidelity, we propose a Dual-Path Self-Corrective Guidance (DSG) strategy. Inspired by CFG (Ho & Salimans, 2021), DSG introduces two concurrent denoising pathways during inference: a non-guided path that relies on the model’s internal priors and yields high-fidelity but uncontrolled outputs, and a guided path conditioned on the warped input trajectory, which enforces camera motion but is prone to artifacts. By utilizing the difference between these two paths at each denoising step, DSG computes a dynamic correction term that softly adjusts the guided path toward the perceptual quality of the non-guided path. This self-corrective mechanism effectively mitigates trajectory-induced degradation while maintaining alignment with the target camera path, thereby improving both the structural integrity and visual quality of the generated video.
@@
Together, these three mechanisms form a cohesive inference-time guidance framework that achieves robust and precise trajectory control while preserving the generalization ability of VDM priors. Our method is fully training-free and plug-and-play, enabling broad applicability across tasks without model retraining. It is also model-agnostic and readily adapts to backbones such as Wan 2.1 (Wan et al., 2025) and SVD (Blattmann et al., 2023), underscoring its general utility. To validate the effectiveness of the proposed method, we conduct comprehensive experiments on multiple tasks and benchmarks. Results demonstrate consistent improvements in trajectory adherence, geometric consistency, and perceptual quality compared to state-of-the-art (SOTA) baselines. Our main contributions are summarized as follows:
@@
  • We introduce a novel, training-free paradigm for leveraging video generative priors in spatial intelligence tasks, enabling precise and stable 3D/4D trajectory control without retraining or fine-tuning.
@@
  • We design a synergistic inference-time guidance framework integrating Intra-Step Recursive Refinement (IRR) and Flow-Gated Latent Fusion (FLF), achieving accurate trajectory adherence while disentangling motion from content.
@@
  • We propose Dual-Path Self-Corrective Guidance (DSG), a self-referential correction mechanism that enhances spatial alignment and perceptual fidelity without auxiliary networks or retraining.
@@
  • We demonstrate, through extensive experiments on diverse datasets and tasks, that our approach achieves state-of-the-art controllability and visual quality, even compared to training-intensive pipelines.
@@ 2 Related Works
We review prior work in three areas most relevant to ours: 3D static scene generation, 4D trajectory-controlled video generation, and guidance strategies for generative models.
3D Static Scene Generation. Recent advances in 3D reconstruction (Mildenhall et al., 2020; Kerbl et al., 2023; Song et al., 2024; Gao et al., 2024; Yu et al., 2024d; Müller et al., 2022; Yao et al., 2018) and object-level generation (Poole et al., 2023; Wei et al., 2024; Xiang et al., 2025; Kwak et al., 2024) have achieved strong results, but they lack scene-level priors and struggle with geometric consistency, limiting their scalability. VDM (Blattmann et al., 2023; Wan et al., 2025; Kong et al., 2024) offer richer priors and have thus become a foundation for scene generation. Approaches vary: some decode 3D scenes from a single image via latent traversal (e.g., Wonderland (Liang et al., 2025)), others fine-tune VDMs with depth-based warped inputs, (e.g., ViewCrafter(Yu et al., 2024c) and See3D (Ma et al., 2025a)), and some methods such as MotionCtrl (Wang et al., 2024c) and TrajectoryAttention (Xiao et al., 2025) embed camera parameters to guide view synthesis . More efficient, training-free strategies such as NVS-Solver (You et al., 2025) and ViewExtrapolator (Liu et al., 2024) warp input views and modulate frozen models during sampling. Fine-tuning offers controllability but risks eroding pretrained priors and incurs high cost, while training-free methods retain priors and efficiency but must address geometric coherence. Our work follows the training-free direction, aiming to ensure both view consistency and controllability.
Trajectory-Controlled Dynamic Video Generation. Video synthesis with controllable camera motion generally follows two main paradigms. The first is fine-tuning-based (Ma et al., 2025b; Mou et al., 2024; Yu et al., 2024b; Wang et al., 2024c) , where lightweight adapters $a_{\phi}$ (e.g., LoRA (Hu et al., 2022), ControlNet (Zhang et al., 2023)) are trained on video–trajectory pairs to optimize the standard diffusion (Ho et al., 2020) loss. Examples include ReCamMaster (Bai et al., 2025a), which retrains on 136K annotated videos, TrajectoryCrafter (Yu et al., 2025) with dual-stream conditioning on source videos and 3D point-cloud renders, GCD (Van Hoorick et al., 2024) using synthetic multi-view videos, and DaS (Gu et al., 2025) incorporating 3D tracking signals into a pre-trained diffusion model for multi-type motion control. The second paradigm is warp-and-repaint, which projects source frames with depth to target poses and then repaints the occluded regions (Ma et al., 2025b; Liu et al., 2025; Huang et al., 2025; Tian et al., 2025). While flexible, this approach remains vulnerable to noisy warps that cause flicker or distortion—as in training-free baselines such as NVS-Solver (You et al., 2025), which modulates a frozen video diffusion model using warped views at inference time. In contrast, our method applies a more powerful inference-time guidance mechanism that extracts trajectory cues and directly steers the diffusion process, achieving precise motion-consistency control without any training.
@@
Guidance and Control for Generative Models. A central challenge in diffusion models is how to steer the generative process toward desired outputs. Guidance strategies address this by modifying the sampling trajectory to better satisfy conditioning signals. The most common is Classifier-Free Guidance (CFG) (Ho & Salimans, 2021), which biases generation toward a target condition by blending conditional and unconditional score predictions. While effective, high guidance weights can cause distortions. More advanced methods, such as Auto-Guidance and STG (Karras et al., 2024; Hyung et al., 2025; Xu et al., 2023), use an auxiliary model to anticipate and avoid failure modes, whereas Z-sampling (Bai et al., 2025c) alternates denoising and inversion to refine results mid-generation. For 3D and 4D synthesis, enforcing viewpoint consistency often follows the fine-tuning and warp-and-repaint strategies discussed above. However, the latter remains sensitive to noisy warps (Cai et al., 2024; Wang et al., 2024a), leading to flicker and geometric distortions. To this end, we propose Dual-Path self-corrective guidance, which derives a correction signal from the difference between guided and unguided predictions at each step, enhancing trajectory adherence and stability without retraining or per-scene tuning.
@@ 3 Proposed Methods
We address the challenge of balancing controllability, visual fidelity, and generalization when applying video diffusion models (VDMs) to 3D/4D tasks. Our solution is a training-free framework for trajectory-controlled video generation. At its core is an inference-time guidance strategy that steers a pre-trained diffusion model along user-defined trajectories while preserving its intrinsic priors and generative quality. As shown in Fig. 2, the framework integrates three complementary components. Intra-Step Recursive Refinement (IRR) injects trajectory guidance from observed regions at each denoising step, ensuring consistent control throughout sampling (see Sec. 3.2). Flow-Gated Latent Fusion (FLF) refines trajectory injection by decoupling motion and appearance features in the latent space, which prevents content drift and preserves fidelity (see Sec. 3.3). Dual-Path Self-Corrective Guidance (DSG) enhances stability by comparing guided and unguided predictions, using their difference as a corrective signal to suppress artifacts from noisy priors (see Sec. 3.4). Together, these modules unlock the model’s latent 3D/4D awareness and enable fine-grained trajectory control without retraining.

### v2

@@ Abstract
Recent video diffusion models show immense potential for spatial intelligence tasks due to their rich world priors, but this is undermined by limited controllability, poor spatial-temporal consistency, and entangled scene-camera dynamics. Existing solutions, such as model fine-tuning and warping-based repainting, struggle with scalability, generalization, and robustness against artifacts. To address this, we propose WorldForge, a training-free, inference-time framework composed of three tightly coupled modules. 1) Intra-Step Recursive Refinement injects fine-grained trajectory guidance at denoising steps through a recursive correction loop, ensuring motion remains aligned with the target path. 2) Flow-Gated Latent Fusion leverages optical flow similarity to decouple motion from appearance in the latent space and selectively inject trajectory guidance into motion-related channels. 3) Dual-Path Self-Corrective Guidance compares guided and unguided denoising paths to adaptively correct trajectory drift caused by noisy or misaligned structural signals. Together, these components inject fine-grained, trajectory-aligned guidance without training, achieving both accurate motion control and photorealistic content generation. Our framework is plug-and-play and model-agnostic, enabling broad applicability across various 3D/4D tasks. Extensive experiments demonstrate that our method achieves state-of-the-art performance in trajectory adherence, geometric consistency, and perceptual quality, outperforming both training-intensive and inference-only baselines.
Figure 1: We present WorldForge, a fully training-free framework leveraging a pre-trained video diffusion model for various 3D/4D tasks, such as monocular 3D scene generation (left) and dynamic 4D scene re-rendering (right), enabling precise camera trajectory control and high-quality outputs.
@@ 1 Introduction
Recent video diffusion models (VDMs) (Blattmann et al., 2023; Wan et al., 2025; Yang et al., 2024; Google DeepMind, 2025) have significantly advanced spatial intelligence (Cao et al., 2025) tasks like 3D/4D understanding (Bahmani et al., 2025a; b), reconstruction (Wang et al., 2025a; Wu et al., 2025; Shi et al., 2024), and generation (Yu et al., 2024c; 2025). Trained on vast video datasets, these models encode rich spatiotemporal priors, enabling realistic spatial transformations for applications like novel view synthesis (You et al., 2025; Xiao et al., 2025), panoramic video (Wang et al., 2024b; Ma et al., 2024a), and dynamic scene reconstruction (Bai et al., 2025a; Yu et al., 2025; Van Hoorick et al., 2024). Furthermore, VDMs are increasingly used to build “world models” (Bar et al., 2025; Duan et al., 2025; Bruce et al., 2024), which are structured internal representations that support predictive reasoning in embodied AI.
@@
Despite their strong priors, VDMs face fundamental limitations, including limited controllability, spatial–temporal consistency, and geometric fidelity, particularly when applied to 3D or 4D tasks (Wang et al., 2024c; He et al., 2024; Ling et al., 2024; Xing et al., 2024). They struggle to follow precise motion constraints, such as a 6-DoF camera trajectory (Hu, 2024; Ma et al., 2024b), which undermines spatial consistency in tasks such as novel view synthesis and trajectory control. These models also entangle scene and camera motion, causing unintended object deformations and scene instability when viewpoint changes are desired (Yu et al., 2024c; Liu et al., 2024). These limitations hinder their use in applications requiring structured spatial reasoning or controllable generation.
@@
To handle these issues, prior works (Jeong et al., 2025; Ren et al., 2025; Yu et al., 2025; Zhang et al., 2025) have pursued two main directions. The first, fine-tuning on motion-conditioned data (Bai et al., 2025a; Xiao et al., 2024; Bai et al., 2025b), can improve control but is computationally costly, generalizes poorly, and risks degrading pretrained priors. The second, a ”warping-and-repainting” strategy (Ma et al., 2025b; Liu et al., 2025; Ma et al., 2025a; You et al., 2025), re-projects frames along a new camera path and uses a generative model to fill occlusions. Although this approach is more flexible, it lacks robustness because pretrained models handle warped, out-of-distribution (OOD) (Yu et al., 2024a) inputs poorly, often producing artifacts and fragmented geometry. Moreover, a bias toward dynamic training data can cause hallucinated motion in static scenes. Consequently, balancing fine-grained controllability with generation quality and generalization remains a challenging open problem.
@@
To address this challenge, we aim to inject precise control into VDMs while preserving their valuable priors. For this purpose, we propose a general inference-time guidance paradigm that leverages the rich priors of VDMs in spatial intelligence tasks, such as geometry-aware 3D scene generation and video trajectory control. Our method uses a warping-and-repainting pipeline, in which input frames are warped along a reference trajectory and then used as conditional inputs in the repainting stage. Building on this, we develop a unified, training-free framework composed of three complementary mechanisms, each designed to address a specific challenge in trajectory-controlled generation.
@@
First, to ensure the generated motion follows the target trajectory derived from depth-based rendering (Wang et al., 2025b; Piccinelli et al., 2024), we introduce Intra-Step Recursive Refinement (IRR). It embeds a micro-scale predict–correct loop within each denoising step: before the next timestep, predicted content in observed regions is replaced with the corresponding ground-truth (GT) observations. This incremental correction allows trajectory control signals to be injected at every step, enabling fine-grained aligned with the target trajectory.
@@
Second, we observe that different channels of the VAE-encoded (Kingma & Welling, 2013; Foti et al., 2022) latent representation encode different information, with some channels specializing in appearance and others in motion. Directly overwriting all channels when injecting trajectory signals can inadvertently degrade visual details. To address this, we propose Flow-Gated Latent Fusion (FLF), which leverages optical-flow similarity to selectively inject trajectory information into channels highly correlated with motion, while leaving appearance-relevant channels unmodified. This process effectively decouples appearance from motion, allowing for precise viewpoint manipulation while preserving content fidelity.
@@
Finally, while warping-based rendering effectively enforces user-defined trajectories, it inevitably introduces noise and visual artifacts stemming from imperfect depth, occlusions, or misalignments (You et al., 2025). To balance control with quality, we propose Dual-Path Self-Corrective Guidance (DSG). Inspired by CFG (Ho & Salimans, 2021), DSG uses two parallel denoising paths during inference: A non-guided path that relies on the model’s priors to produce high-fidelity but uncontrolled results, and a guided path that follows the warped trajectory, ensuring camera control but risking artifacts. At each step, DSG computes the difference between these paths to create a dynamic correction term. This term adjusts the guided path toward the higher perceptual quality of the non-guided path. This self-corrective mechanism mitigates artifacts from the warped trajectory while maintaining camera control, improving the video’s overall structure and visual quality.
@@
Together, these three mechanisms form a cohesive inference-time guidance framework for robust and precise trajectory control while preserving VDM priors. Our method is training-free and plug-and-play, enabling broad applicability across tasks without model retraining. It is also model-agnostic and readily adapts to backbones such as Wan 2.1 (Wan et al., 2025) and SVD (Blattmann et al., 2023). Comprehensive experiments on multiple tasks and benchmarks confirm that our approach achieves state-of-the-art (SOTA) results, improving trajectory adherence, geometric consistency, and perceptual quality over leading baselines. Our main contributions are:
@@
  • A novel, training-free paradigm for leveraging VDM priors in spatial intelligence tasks, enabling precise and stable 3D/4D trajectory control without retraining or fine-tuning.
@@
  • A synergistic inference-time guidance framework integrating Intra-Step Recursive Refinement (IRR) and Flow-Gated Latent Fusion (FLF), achieving accurate trajectory adherence while disentangling motion from content.
@@
  • Dual-Path Self-Corrective Guidance (DSG), a self-referential correction mechanism that enhances spatial alignment and perceptual fidelity without auxiliary networks or retraining.
@@
  • Extensive experiments on diverse datasets and tasks show our approach achieves SOTA controllability and visual quality, even compared to training-intensive pipelines.
@@ 2 Related Works
We review prior work in three relevant areas: 3D static scene generation, 4D trajectory-controlled video generation, and guidance strategies for generative models.
3D Static Scene Generation. While 3D reconstruction (Mildenhall et al., 2020; Kerbl et al., 2023; Song et al., 2024; Gao et al., 2024; Yu et al., 2024d; Müller et al., 2022; Yao et al., 2018) and object generation (Poole et al., 2023; Wei et al., 2024; Xiang et al., 2025; Kwak et al., 2024) are advanced, they often lack scene-level priors. VDM (Blattmann et al., 2023; Wan et al., 2025; Kong et al., 2024) provide these priors and are leveraged by decoding scenes from images (Liang et al., 2025), fine-tuning on warped inputs (Yu et al., 2024c; Ma et al., 2025a), or embedding camera parameters (Wang et al., 2024c; Xiao et al., 2025). Unlike costly fine-tuning which can corrupt priors, training-free strategies (You et al., 2025; Liu et al., 2024) preserve them but must ensure geometric coherence. Our work takes this training-free approach to enhance both consistency and control.
Trajectory-Controlled Dynamic Video Generation. One paradigm for controllable video synthesis is to fine-tune lightweight adapters (Ma et al., 2025b; Mou et al., 2024; Yu et al., 2024b; Wang et al., 2024c) like LoRA (Hu et al., 2022) or ControlNet (Zhang et al., 2023) on video-trajectory data, conditioning on diverse inputs (Bai et al., 2025a; Yu et al., 2025; Van Hoorick et al., 2024; Gu et al., 2025). Another is the warp-and-repaint strategy (Ma et al., 2025b; Liu et al., 2025; Huang et al., 2025; Tian et al., 2025), which projects and inpaints frames but is prone to artifacts from noisy warps (You et al., 2025). Our work uses inference-time guidance to directly steer the diffusion process for precise motion control.
@@
Guidance and Control for Generative Models. Guidance strategies steer diffusion models toward desired outputs. While Classifier-Free Guidance (CFG) (Ho & Salimans, 2021) is common, high weights can cause artifacts. More advanced techniques use auxiliary models (Karras et al., 2024; Hyung et al., 2025; Xu et al., 2023) or iterative refinement (Bai et al., 2025c) to improve sampling. In 3D/4D synthesis, guidance is used to enforce viewpoint consistency, but warp-based methods often suffer from noise-induced artifacts (Cai et al., 2024; Wang et al., 2024a). To address this, we propose DSG. It derives a correction signal from the difference between guided and unguided predictions at each step, enhancing trajectory adherence and stability without retraining.
@@ 3 Proposed Methods
We propose an inference-time guidance strategy to balance controllability with visual fidelity for VDMs in 3D/4D tasks. Our method is a training-free framework that steers a pretrained model along a user-defined trajectory while preserving its generative priors. As shown in Fig. 2, our framework has three key components. First, Intra-Step Recursive Refinement (IRR) injects trajectory guidance from observed regions at each denoising step for consistent control (Sec. 3.2). Second, Flow-Gated Latent Fusion (FLF) decouples motion from appearance in the latent space to prevent content drift and preserve fidelity (Sec. 3.3). Finally, Dual-Path Self-Corrective Guidance (DSG) uses the difference between guided and unguided paths as a corrective signal to suppress artifacts and improve stability (Sec. 3.4). Together, these modules enable fine-grained trajectory control and unlock the model’s latent 3D/4D awareness without any retraining.